9.1.Find the force of attraction betweenthenucleus of a hydrogen


atom and an electron. The radius of the hydrogen atom isO.5x10-8cm


and the charge of the nucleus is equal in magnitude and opposite in


sign to that of the electron.

1
Expert's answer
2022-01-27T13:16:25-0500

Solution

Charge of electron

Q"=1.6\times10^{-19}C"

Radius

R=O.5x10-8 cm

Then electrostatic force between electron and nucleus

"F=\frac{KQQ}{R^2}\\=\frac{9\times10^9\times1.6\times10^{-19}\times 1.6\times10^{-19} }{( 0.5\times10^{-10}) 
 ^2}\\=92.16\times10^{-9}N."